To debug a problem with the new c1psl (later elog), we needed a Supermicro EPICS server that was using the shared EPICS/modbus/asyn binaries rather than a local install. Of those available in the lab (c1iscaux, c1vac, c1susaux being the others), this was the only one which uses the shared install. So I
At which point Jon reset the software end, I restored the slow bias voltage and re-enabled the local damping. The optic seems to have damped okay. The Oplev spot is back in ~center of the QPD and the green beam can be locked to a TEM00 mode (so the alignment is okay - the IR beam is unavailable while c1psl issues are being sorted but I judge that things are back to the nominal state now).